A small open economy is described by the following set of equations:

C = 300 + 0.6(Y − T)

I = 700 − 80r

NX = 200 − 50ε

G = T = 500 (Balanced Budget)

(M/P)^d = Y − 200r

M = 3, 000

P = 3

r ∗ = 5

(a) Derive and graph the IS∗ and LM∗ curves.

(b) Calculate the equilibrium exchange rate, income and net exports.

(c) Assume a floating exchange rate. Calculate what happens to the exchange rate, income, net exports, and the money supply if the government increases its spending by 200. Use a graph to explain what you find.

(d) Now assume a fixed exchange rate. Calculate what happens to the exchange rate, income, net exports, and the money supply if the government increases its spending by 200. Use a graph to explain what you find.
Answer #1

January 2013 01/01/2013 Week 01 - Tuesday 07 Thought of the Day There is nothing more frightful than ignorance in action 8 am

Thought of the Day Take time by the forelock 8 am 9 Nx = 200-50 (4) Ny = 200 - 200 NX-o. Lequilibrum net experts at =4 equati

Thought of the Day Difficulty is a severe instructor 8 am in the The LM curve is unchanged because there is no Change supply